# A Low Duty Cycle Pulsed UV Technique for Spectroscopy of Aluminum Monochloride

### v1

@@ 2.3 Characterization of Long-Term Behaviour and Further Improvements
Using our technique, the measurements in Fig. 3 were conducted over a two-day period during which no degradation of the BBO crystal was observed. The primary reason for the long duration of the experimental run is the low repetition rate of $\approx 1$ Hz and the high number of averages per data point. The experiment is run at such a slow rate to keep the heat load from the ablation laser on the cryogenic cell at a minimum, as is typically done in CBGB experiments. High-number averaging is necessary to compensate for variations in the molecular yield when the ablation laser rasters over the target.
@@
To characterize the long-term behaviour of our method, we monitored the output intensity over a long period of time, as shown in Fig. 4. The average intensity variations between frequency points in the scans were mostly below $10\%$ for at least 6 hours without any cavity alignment optimization. We attribute these fluctuations primarily to set point drifts in the SHG cavity stabilization circuitry. While these fluctuations introduce no systematic frequency offsets in the experiment, improved power stability could be achieved by incorporating an acousto-optic modulator at the output of the cavity for intensity stabilization [72].
@@
We note that the SHG cavity can drift in and out of resonance outside the experimental sequence window when the piezo is held at a constant voltage. This effect leads to unwanted threshold trigger events that switch on the cavity feedback loop. A simple way to avoid this issue is to only arm the threshold trigger at the beginning of the applied voltage ramp. Moreover, if it is desirable to avoid that the cavity drifts into resonance at any time, an easy mitigation technique would be to monitor the output intensity outside the sequence window and use another threshold trigger in combination with an inverted PI control to deliberately push the cavity out of resonance. For the data presented in this work, we didn’t restrict the trigger window and relied on occasional manual adjustments of the offset voltage of the ramp to circumvent the issue since we did not observe any degradation due to the sporadically resonant cavity.
@@
To increase the UV output power, e.g. for laser cooling purposes, in our home built system, we also explored using a CLBO crystal (Conex Optics Inc.) for the SHG conversion. This choice of crystal is motivated by the fact that CLBO has an overall higher conversion efficiency than BBO at our fundamental wavelength of 523 nm, which is mainly attributed to the reduced walk-off angle of CLBO [27, 73]. Overall, the cavity UV output increased by an order of magnitude using the CLBO. However, when we used the cavity in continuous mode, i.e. before the development of the pulsed technique, we observed a severe degradation of the crystal performance over the course of minutes. The cavity output power could be brought back to a maximum by translating the crystal, which in turn moves the focus and the entry and exit points of the fundamental cavity mode to a different spot. Translating the focus back to the original point resulted in the low output power again. This behaviour was observed both with BBO and CLBO crystals, both being at room temperature, and prompted the development of the low-duty cycle technique. In addition, in the case of CLBO, we observed fractures in the crystal that developed over the course of several days of operation, as shown in Fig. 5. We attribute these damages to the high sensitivity of CLBO to the presence of residual moisture and dust in our cavity enclosure, which is not perfectly sealed. The building of cracks in CLBO due to water is a known fact and has been studied previously [74, 75]. However, the operation of the cavity in continuous mode seemingly accelerated this effect. On the other hand, taking advantage of the pulsed technique, we were able to operate the SHG cavity with a new CLBO crystal for several weeks without any notable crystal damage.
@@
Finally, we note that any technique that runs an experimental setup with a reduced duty cycle exhibits pulsed heat loads, resulting in a non-equilibrated state of the setup. In this case of an SHG cavity, the absorption of the light in the crystal leads to what is known as self-heating, which has been observed with both BBO and CLBO crystals, especially in high-power UV systems [32, 11, 9]. As a consequence, the phase matching condition is slightly altered due to different net heat loads between the configuration when the user aligns the cavity setup for optimal output power and when the setup is used for an experimental run. This change leads to a reduction of the conversion efficiency and can be mitigated in different ways. In one approach, users align and optimize the SHG cavities in a configuration that mimics the experimental run in terms of, for instance, repetition rates and used laser powers as close as possible to operate under the same heat loads. In a second approach, users intentionally lower the temperature of their crystals, slightly deviating from the optimal phase matching condition, as has been shown in previous works [11, 9, 29]. Yet another way is to maintain the crystal at an elevated temperature such that the relative temperature increase due to self-heating is small compared to the crystal temperature. The latter approach is best applied to crystals that use angle, instead of temperature, crystal phase-matching, which is the case for BBO and CLBO around 500 nm. Heating the crystal is desirable in any case since the detrimental effects of moisture can be reduced. CLBO, for instance, has been successfully operated without degradation at an elevated temperature of 150 deg C and produced stable Watt-level UV output power, as described in Ref. [27]. We note that a major difference of that experiment is the use of commercial cavities made of stainless steel enclosures, which are well sealed and can readily be kept clean and dry. Nevertheless, especially at such high UV output powers, most optics to guide or manipulate the laser light after the cavity is susceptible to UV-induced damage. Our pulsed approach will minimize the impact of the UV light and prolong the lifetime of the SHG cavity and the following optics as well.
@@ 3 Conclusion
In this work, we have developed and tested a technique to use a deep-UV SHG laser system while reducing UV-induced damage on optics to a bare minimum. Our approach to reduce the duty cycle of the used UV light is ideal for any pulsed experiments but can also be applied in CW experiments to bridge, often unavoidable, dead times and reduce the UV operation time as much as possible. The frequency accuracy of the laser system should not be affected when using this method since the ramping of the piezo-actuated mirror only affects the output intensity and the frequency stability remains with the fundamental laser. The long-term stability of our setup is sufficient for typical time scales of atomic and molecular experiments. While our technique is particularly well-suited for protecting non-linear crystals in resonant-enhancement cavities where the circulating intensity of the fundamental laser light is large, any subsequent optical components in the experiment, such as mirrors and optical fibers, experience less degradation as well due to the reduced average power of the used UV light. We applied this technique to measure the isotope shift of the electronic $X^{1}\Sigma$-$A^{1}\Pi$ transition in AlCl and found excellent agreement with ab initio calculations.

### v2

@@ 2.3 Characterization of Long-Term Behaviour and Further Improvements
Using our technique, the measurements in Fig. 4 were conducted over a two-day period. The primary reason for the long duration of the experimental run is the low repetition rate of $\approx 1$ Hz and the high number of averages per data point. The experiment is run at such a slow rate to keep the heat load from the ablation laser on the cryogenic cell at a minimum, as is typically done in CBGB experiments. High-number averaging is necessary to compensate for variations in the molecular yield when the ablation laser rasters over the target.
@@
To characterize the long-term behaviour of our method, we monitored the output intensity over a long period of time, as shown in Fig. 5. The average intensity variations between frequency points in the scans were mostly below $10\%$ for at least 6 hours without any cavity alignment optimization. We attribute these fluctuations primarily to set point drifts in the SHG cavity stabilization circuitry and the temperature sensitivity of polarizing optical components. Overall, the drifts are comparable to what has been observed in other systems [93, 104]. However, while these fluctuations introduce no systematic frequency offsets in the experiment, an improved power stability could be achieved by incorporating an acousto-optic modulator at the output of the cavity for intensity stabilization [104].
@@
We note that the SHG cavity can drift in and out of resonance outside the experimental sequence window when the piezo is held at a constant voltage. This effect leads to unwanted threshold trigger events that switch on the cavity feedback loop. A simple way to avoid this issue is to only arm the threshold trigger at the beginning of the applied voltage ramp. Moreover, if it is desirable to avoid that the cavity drifts into resonance at any time, an easy mitigation technique would be to monitor the output intensity outside the sequence window and use another threshold trigger in combination with an inverted PI control to deliberately push the cavity out of resonance. For the data presented in this work, we didn’t restrict the trigger window and relied on occasional manual adjustments of the offset voltage of the ramp which was sufficient to run the experiment.
@@
To increase the UV output power, e.g. for laser cooling purposes, in our home built system, we also explored using a CLBO crystal (Conex Optics Inc.) for the SHG conversion. This choice of crystal is motivated by the fact that CLBO has an overall higher conversion efficiency than BBO at our fundamental wavelength of 523 nm, which is mainly attributed to the reduced walk-off angle of CLBO [28, 105]. Overall, the cavity UV output increased by an order of magnitude using the CLBO. However, when we used the cavity in continuous mode, i.e. before the development of the pulsed technique, we observed a severe degradation of the crystal performance over the course of minutes. The cavity output power could be brought back to a maximum by translating the crystal, which in turn moves the focus and the entry and exit points of the fundamental cavity mode to a different spot. Translating the focus back to the original point resulted in the low output power again. This behaviour was observed both with the BBO and CLBO crystals. In addition, in the case of CLBO, we observed fractures in the crystal that developed over the course of several days of operation, as shown in Fig. 2.
We attribute these damages to operating the non-linear crystals at room temperature and to their high sensitivity to the presence of humidity in the air near the crystal. For both crystal types, the negative effects of elevated humidity have been studied extensively. In the case of CLBO, it has been observed that in particular humidity levels higher than 40% are detrimental to the crystal’s performance [106, 107]. While at lower levels, water still gets incorporated in the crystal structure but does not degrade the crystal performance, at levels above 40% water reacts with the CLBO itself to form polycrystals at the surface that have a different crystal phase [108]. Moreover, exposing a hydrated crystal with deep-UV light at 266 nm leads to the formation of OH groups, which in turn leads to damaging photo-assisted chemical reactions on the crystal surface [109]. Finally, studies have shown that water molecules preferably penetrate the facets parallel to the CLBO’s c-axis, leading to cracks due to hydrolysis within days when exposed to humidity levels higher than 40% [106, 107]. For BBO, a two-photon absorption process in the deep UV is suggested to be responsible for creating color centers. These defects lead to localized heating due to absorption at longer wavelengths, which reduces the conversion efficiency due to localized changes of the phase-matching condition [110, 111, 112, 113, 114, 115, 116, 117].
Consequently, it is recommended to keep such non-linear crystals in a dry environment and use high-quality crystals since internal defects are known to form cracks more easily, at least for CLBO [106]. More importantly, the hydration and degradation of the non-linear crystals can be drastically reduced when keeping the crystal at high temperatures. For CLBO, operating temperatures of 150 C have shown to mitigate the degradation effects and to allow for continuous operation of a high-power SHG cavity [118, 28]. Equivalently, for BBO at operating temperature of 200 C, a 3.5-fold reduction of the two-photon absorption coefficient has been observed, allowing for continuous operation of a SHG cavity [117].
While our enclosure of our SHG cavity provides a clean and relatively dry environment during the experiment, any manual alignment of the cavity happens when the enclosure is open which exposes the crystals to the ambient laboratory air with humidity levels often much higher than 40%. This severely limits the ability to improve the cavity alignment to improve the UV output power and we assume that this external water exposure is the most likely cause of the cracked CLBO that we observed. In addition, we observed that the operation of the SHG cavity in continuous mode seemingly accelerated the degradation effect of the crystal. On the other hand, our developed pulsed technique allowed us to operate the homebuilt SHG cavity for a prolonged time even at room temperature without any notable crystal damage. It is likely that a side effect of reducing the UV-duty cycle is that any detrimental photo-assisted effects and damages in the crystal are reduced.
In the future, we plan to add a temperature control to maintain the non-linear crystal at high temperatures above 130  C at all times to mitigate these challenges and develop a laser system at 260 nm with Watt-level output powers. Here, reducing the duty cycle of the UV light will be paramount to protect any involved optical elements after the SHG cavity. However, while we expect that the elevated temperature of the crystal removes the hydration and corresponding degradation issues, there are several aspects that need to be considered when employing the pulsed technique in a high power setting.
@@
Any technique that runs an experimental setup with a reduced duty cycle exhibits pulsed heat loads, resulting in a non-equilibrated state of the setup. In this case of a SHG cavity, the absorption of the light in the crystal leads to what is known as self-heating, which has been observed with both BBO and CLBO crystals, especially in high-power UV systems [33, 12, 9]. As a consequence, the phase matching condition is slightly altered due to different net heat loads between the configuration when the user aligns the cavity setup for optimal output power and when the setup is used for an experimental run. This change leads to a reduction of the conversion efficiency and can be mitigated in different ways. In one approach, users align and optimize the SHG cavities in a configuration that mimics the experimental run in terms of, for instance, repetition rates and used laser powers as close as possible to operate under the same heat loads. In a second approach, users intentionally lower the temperature of their crystals, slightly deviating from the optimal phase matching condition, as has been shown in previous works [12, 9, 30]. Yet another way, which we plan to employ, is to maintain the crystal at an elevated temperature such that the relative temperature increase due to self-heating is small compared to the crystal temperature. The latter approach is best applied to crystals that use angle, instead of temperature, crystal phase-matching, which is the case for BBO and CLBO around 500 nm.
Finally, one expects that the thermal shock of the laser absorption deforms the lattice of the nonlinear crystal, eventually leading to a decrease of the conversion efficiency. This effect, however, has been studied in detail in CLBO and it has been shown that when operating at temperature above 130 deg C long-term operation is possible since the thermal stress is minimized due to activated phonon vibration [118].
@@ 3 Conclusion
In this work, we have developed and tested a technique to reduce the duty cycle of a deep-UV SHG laser system. Our approach is ideal for any pulsed experiments but can also be applied in CW experiments to bridge, often unavoidable, dead times and reduce the UV operation time as much as possible. The frequency accuracy of the laser system should not be affected when using this method since the ramping of the piezo-actuated mirror only affects the output intensity and the frequency stability remains with the fundamental laser. The long-term stability of our setup is sufficient for typical time scales of atomic and molecular experiments. We applied this technique to measure the isotope shift of the electronic $X^{1}\Sigma$-$A^{1}\Pi$ transition in AlCl and found excellent agreement with ab initio calculations.
[…]